OpenStudy (anonymous):

what is the solution set of the quadratic equation 9x^2 + 7x + 2=0?

OpenStudy (anonymous):

it has imaginory roots..

OpenStudy (anonymous):

thx and how do you figure that out

OpenStudy (anonymous):

b^2 - 4ac is negative

OpenStudy (anonymous):

the discriminant 7^2-4*9*2=49-72<0...

OpenStudy (anonymous):

for a given quadratic equation, ax^2+bx+c find the determinant D=sqrt(b^2-4ac) if D<0 then the equation has a imaginary set of roots.

OpenStudy (anonymous):

sorry D is discriminant ! :|